### v2

@@ 2.7 Synthetic data generation
Next, we used the fine-tuned model from Section 2.6 to generate motion mode informed synthetic light curves (Figure 10B, C). To do this, we first took a held-out test set used to evaluate the fine-tuned model from Section 2.6. Next we obtained reference curves for the latent space by using light curves with high confidence predictions from that model (>95% likelihood). Finally, we generated around the neighbourhood of those references by first applying noise to those latent activations and then decoding into light curves as in Figure 10A. From this, we saw that in each motion class, we could in essence query an inference dataset to produce more curves of a similar attitude motion mode. Considering that producing these fine-tuning data is time and computationally expensive, we present this and the methodology as something of potentially high utility to the SDA field.
@@ 3.1 Data Sources and Preparation
Table 4: Data modalities and their sources
Data Modality Source Notes n
Light curves MMT-9 Real obs. 227,697
(10,600 sats)
CASSANDRA Simulated. 800
GMV GRIAL Simulated. 22,009
The MMT-9 light curve dataset was downloaded as text files from https://www.sao.ru/lynx/karpov/satellites/ and processed with custom scripts using the BeautifulSoup Python library (v4.11.1). The light curves varied in length, and were therefore resampled to a uniform size of $n_{samples},128$. Visual inspection showed this length to not noticeably degrade light curves. Similarly, the anomaly detection (CASSANDRA) and motion prediction (GRIAL) fine-tuning datasets were processed to $n_{samples}$$,128,n_{labels}$, where there were two labels for anomaly detection and six for motion prediction. For details on dataset sizes, please see Table 4, which described the datasets produced by the CASSANDRA simulator (anomaly detection fine-tuning), and GRIAL simulator (motion prediction fine-tuning).
@@ 3.2 Model Architecture
Our Perceiver-VAE model archiecture is outlined in Equation 1, and we briefly describe these steps below.
@@ 3.3 Pre-training strategy
We pre-trained our Perceiver-VAE model using a multi-task self-supervised learning approach. For this we incorporated three objectives: reconstruction, masking, and forecasting. The reconstruction task trained the model to encode and decode whole light curves, establishing the foundations for latent space representation. At the same time, the masking task required the model to predict randomly masked segments of input sequences, encouraging our model to learn contextual relationships within the light curve. Finally, the forecasting task trained the model to predict future sequence values, promoting the capture of temporal dynamics and enabling forecasting at inference time. We combined these learning objectives by summing individual mean squared error loss components for each task, combined with a KL divergence term (weight $\alpha$0.001) to regularise the latent space (Equation 3. This combined otpimisation enabled the model to learn rich representations of light curve characteristics while maintaining a structured latent space suitable for generation and interpolation.
@@
Pre-training was conducted with a batch size of 32 and an initial learning rate of $1e^{-3}$, with gradient clipping at 0.5 to prevent exploding gradients. We implemented a learning rate scheduler that reduced the rate by a factor of 0.5 when validation loss plateaued for 5 consecutive epochs. We pre-trained for 50-200 epochs, implementing early stopping with a patience of 50 epochs to ensure proper convergence while preventing overfitting.
@@ 3.4 Fine-tuning
For fine-tuning, we provide an additional constraint on the model. In addition to the Variational autoencoding task (Equation 3), the model also minimises binary cross entropy loss (Equation 9, or categorical cross entropy loss in the case of the multi-class motion prediciton problem). This results in an overall loss term of:
$\mathcal{L}_{\text{fine-tune}}=\mathcal{L}_{\text{total}}+\mathcal{L}_{\text{class}},$ (8)
where $\text{where }\mathcal{L}_{\text{class}}=-\frac{1}{m}\sum_{i=1}^{m}[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})].$ (9)
Where the classification loss employs binary cross-entropy across $m$ samples, measuring the discrepancy between true labels $y_{i}$ and predicted probabilities $\hat{y}_{i}$.
Equation 8 ensures that the model maintains the ability to reconstruct the fine-tuning data whilst making a classification based on the latent space. Note that some fine-tuning approaches remove decoding from the optimisation completely, just optimising the cross entropy loss between the latent variables and the class labels. However, in our case, we use this multi-task fine-tuning setup to reduce the risk of catastrophic forgetting of the model’s ability to reconstruct, forecast and generate synthetic data.
3.4.1 Anomaly Detection Fine-tuning:
During fine-tuning for anomaly detection, we optimised the combined loss function (Equation 8) using Adam with a learning rate of $1e^{-4}$. This approach allowed the model to learn discriminative features for anomaly detection while preserving reconstruction capabilities. We implemented stratified data splitting to handle class imbalance and evaluated performance using both classification accuracy and ROC AUC.
3.5 Motion Prediction Fine-tuning:
For motion prediction fine-tuning, we adapted our pre-trained Perceiver-VAE to classify satellite motion patterns from light curves. We again implemented stratified class-balanced sampling to handle the uneven distribution of motion types in the fine-tuning data. To improve generalization, we applied a learning rate scheduler that reduced the rate by half when the validation loss plateaued for 10 epochs.
@@ 3.6 Synthetic Data Generation
Here we use the fine-tuned motion prediction classifier to generate a set of high confidence ’query’ curves. The algorithm first identifies some reference curves above a given confidence threshold (0.9) for a requested motion class e.g., tumbling. Then, it uses these to index the learned latent space of the model, and samples around that latent space by adding Gaussian noise to the highest activating neurons. These perturbed latent vectors are processed through the decoder to reconstruct synthetic light curves that represent variations within the same motion class. To enhance visual quality and reduce high-frequency artifacts, we applied Gaussian smoothing ($\sigma=2.0$) to the reconstructed signals.
@@ 4 Outlook
Our work is the first to demonstrate that a VAE architecture can effectively learn meaningful representations from light curve data and be fine-tuned for multiple downstream tasks including detecting anomalous space object behaviour and motion prediction. We have also shown that our approach scales well to large datasets while maintaining computational efficiency, which is critical considering the rapidly growing amount of Space Object data. Here, we developed a two-stage approach, first through unsupervised pre-training, whereby the model learns rich features which aid reconstruction of real light curves. Then we performed supervised fine-tuning, which leverages those features for both anomaly classification, reaching 88% accuracy with a 0.9 ROC AUC score, and attitude mode motion classification, achieving 82.7% test accuracy, with 0.95 ROC AUC scores. Additionally, we demonstrate that the same architectural framework and approach can simultaneously achieve multiple space object behavioural analysis goals (i.e., anomaly detection, attitude mode classification, and synthetic data generation), provided that the representations learned during pre-training are diverse and rich, and that these are coupled with high-fidelity fine-tuning simulation data. Although in this study we identify several predicted anomalous light curve modes, some present clearly anomalous behaviour (e.g., rapid tumbling), while for others the underlying behaviour is less clear. Systematic cataloguing and analysis of simulation conditions that lead to these various curve morphologies would allow us to verify whether particular erroneous space object dynamics produce these curves (e.g., in Fig 8C, F).
@@
From this work, several key directions for further development emerge. Firstly, development of robust quantitative benchmarks for AI synthetic data. Whilst our approach encourages plausibility in the generated curves through reference-based sampling, systematic evaluation frameworks are needed to validate produced synthetic data beyond visual inspection/sense checks. For example, these frameworks could assess statistical distribution matching between real and AI generated datasets, and evaluate like-for-like comparisons in downstream tasks trained on real data, numerically simulated data, and/or data generated through our approach (i.e., generative AI data). Finally, physics-informed neural networks (PINNs) could be used to quantify physical plausibility. Briefly, these networks combine a data loss term with a physics loss term. For example, for light curves this might be a reflectivity model as outlined in [20]. At inference time, a well-trained PINN will give a good indicator as to whether a synthetically generated light curve is obeying the underlying physics due to the value of the physics loss term.
@@
A substantial challenge direction for further development is multi-modal fusion - i.e., integrating the light curve analysis presented here with other observation types (e.g., hyperspectral data, orbits, radar etc.) Indeed, part of the decision to use a Perceiver based architecture was to readily enable extension to other data modalities, as it is well suited to this through its cross-attention and latent bottleneck design, which could enable multiple modalities to be reduced to a common set of embeddings. However, there are substantial technical challenges in multimodal fusion, including developing an appropriate fusion strategy at different levels of the pre-training/fine-tuning pipeline. One such method to fuse information from other sensors with light curves could include implementation of contrastive learning using paired observations. For example, positive pairs (same space object at different dates) and negative pairs (different objects on the same date) would help develop more discriminative features for space object characterisation. Additionally incorporating metadata prediction tasks as part of the self-supervised framework could help capture characteristic signatures of specific satellites, or debris. For example, a fine-tuned metadata prediction FM could generate a likely light curve given an orbital eccentricity, which could help unpick the relationships between satellite structure/behaviour and orbital path.
Our initial framework provides promising results, and as the amount of data available increases, a basis for further analysis. We suggest foundation Models could become powerful tools for automated space object monitoring and space situational awareness. As orbital populations continue to grow, automated approaches such as ours will become increasingly crucial for maintaining space safety & sustainability.

### v3

@@ 2.7 Synthetic data generation
Next, we used the fine-tuned model from Section 2.6 to generate motion mode informed synthetic light curves (Figure 9B, C). To do this, we first took a held-out test set used to evaluate the fine-tuned model from Section 2.6. Next we obtained reference curves for the latent space by using light curves with high confidence predictions from that model (>95% likelihood). Finally, we generated around the neighbourhood of those references by first applying noise to those latent activations and then decoding into light curves as in Figure 9A. Without real labelled light curves to compare against, validating these synthetic data is challenging. To this end, we provide a quantitative assessment of synthetic light curve quality, via validation of the generated curves against a Gaussian Process (GP) fitted to the real light curve distribution for each motion class (Figure 10). For each class, high-confidence real curves (predicted probability >>0.75) were used to fit a GP, modelling the class distribution as a non-parametric mean and standard deviation at each time step. Synthetic curves were then scored by the fraction of time steps falling within the GP’s 95% confidence interval (see Section 3.5 for full details).
Across all six motion classes, synthetic curves scored highly against the GP confidence band, with five of six classes exceeding 94% of time steps within the 95% confidence interval (Figure 10B). The highest scoring classes were TUMBLING (97.6%) and INERTIAL (97.5%), both exhibiting distinctive and consistent morphologies in real data, providing a tight distributional prior for the GP. The lowest scoring class, YAW (89.4%), reflects greater morphological variability within that class — a consequence of YAW grouping four distinct sub-motions (YAWXC, YAWXS, YAWZC, YAWZS), which introduces distributional heterogeneity. These results demonstrate that the synthetic curves produced by our neighbourhood sampling approach are statistically consistent with real light curve distributions across motion classes, without requiring manual smoothing. From this, we saw that in each motion class, we could in essence query an inference dataset to produce more curves of a similar attitude motion mode. Considering that producing these fine-tuning data is time and computationally expensive, we present this and the methodology as something of potentially high utility to the SDA field.
@@
Figure 10: Gaussian Process validation of synthetic light curves. (A) Overview of the validation procedure: high-confidence real light curves are selected per class, a GP is fitted to the real distribution, and synthetic curves are scored against the resulting confidence band. (B) GP validation panels for all six motion classes. The orange band and dashed line show the GP 95% confidence interval and mean respectively, fitted to real curves (blue). Synthetic curves are shown in red. The score reported in each panel title denotes the fraction of synthetic curve time steps falling within the GP 95% confidence interval, averaged across all synthetic curves for that class. Five of six classes exceed 94%, with YAW showing a lower score (89.4%) consistent with the greater morphological variability introduced by grouping four sub-motions into a single class.
2.8 Model Compression via Magnitude Pruning
To assess parameter efficiency of the large Perceiver-VAE, we applied magnitude pruning separately to the Perceiver layers and the VAE bottleneck. The Perceiver layers tolerate up to 70% weight sparsity with negligible degradation in motion classification accuracy (82.5% vs. 83.0% baseline). In contrast, the VAE bottleneck is considerably more sensitive: 50% sparsity reduces accuracy to 77.2%, a drop of approximately 6 percentage points. This asymmetry suggests that whilst the Perceiver layers contain substantial redundancy at this scale, the bottleneck representations are dense and critical to downstream performance. Re-pretraining a compact architecture informed by these sparsity patterns is a natural next step, discussed further in Section 4. Furthermore, post-training quantisation of the large model to INT8 precision results in negligible loss in motion classification accuracy, whilst reducing model size by 75% (1,390MB to 350MB on disk), offering a practical route to deployment in resource-constrained environments.
@@ 3.1 Data Sources and Preparation
Table 7: Data modalities and their sources. For GMV GRIAL, training sample counts per motion class are: I=Inertial, SA=Safe, SP=Spin, SU=Sun, T=Tumbling, Y=Yaw.
Source Notes n Labels
MMT-9 Real obs. 227,697 -
CASSANDRA Simulated 800 160 anom., 640 normal
GMV GRIAL Simulated 22,009 I:1259, SA:1940, SP:1280, SU:940, T:1333, Y:5131
The MMT-9 light curve dataset was downloaded as text files from https://www.sao.ru/lynx/karpov/satellites/ and processed with custom scripts using the BeautifulSoup Python library (v4.11.1). The light curves varied in length and were therefore resampled to $n_{\text{samples}}=128$ time steps. Whilst the Perceiver architecture does not strictly require uniform input length, uniform resampling simplifies batching and ensures consistent positional encoding across samples. Visual inspection showed this length to not noticeably degrade light curves. Similarly, the anomaly detection (CASSANDRA) and motion prediction (GRIAL) fine-tuning datasets were processed to $(n_{\text{samples}},128,n_{\text{labels}})$, where there were two labels for anomaly detection and six for motion prediction. For details on dataset sizes, please see Table 7, which described the datasets produced by the CASSANDRA simulator (anomaly detection fine-tuning), and GRIAL simulator (motion prediction fine-tuning).
@@ 3.2 Model Architecture
Our Perceiver-VAE model architecture is outlined in Equation 1, and we briefly describe these steps below.
@@
We evaluate three model scales: small (64 latents, $d_{z}$=128, 6.4M parameters), medium (128 latents, $d_{z}$=128, 13.2M parameters), and large (384 latents, $d_{z}$=512, 347.5M parameters), the latter representing the largest scale trainable within our computational budget. We note that magnitude pruning of the large model retains performance with 60–70% of weights removed, suggesting substantial redundancy; we discuss this further in Section 4.
@@ 3.3 Pre-training strategy
We pre-trained our Perceiver-VAE model using a multi-task self-supervised learning approach. For this we incorporated three objectives: reconstruction, masking, and forecasting. The reconstruction task trained the model to encode and decode whole light curves, establishing the foundations for latent space representation. At the same time, the masking task required the model to predict randomly masked segments of input sequences, encouraging our model to learn contextual relationships within the light curve. Finally, the forecasting task trained the model to predict future sequence values, promoting the capture of temporal dynamics and enabling forecasting at inference time.
Each sample is presented to the model once per batch in a single forward pass. The reconstruction, masking, and forecasting outputs are computed simultaneously within this pass, with each task operating on a different view of the same input (full sequence, randomly masked sequence, and future-masked sequence respectively). The KL divergence is therefore computed once per sample from the single latent sample, and the effective batch size equals the nominal batch size throughout pre-training. We combined these learning objectives by summing individual mean squared error loss components for each task, combined with a KL divergence term (weight $\alpha$0.001) to regularise the latent space (Equation 3. This combined optimisation enabled the model to learn rich representations of light curve characteristics while maintaining a structured latent space suitable for generation and interpolation.
@@
Pre-training hyperparameters were selected via Bayesian optimisation using Optuna [1], running 25 trials of 15 epochs each per model scale. The optimal learning rates were $1.41\times 10^{-3}$, $5.56\times 10^{-3}$, and $5.38\times 10^{-4}$ for small, medium, and large respectively, with batch sizes of 16, 32, and 16. All models were trained with gradient clipping at 0.5 [22] to prevent exploding gradients, which we observed empirically during the early training phase where the KL divergence term increases rapidly before stabilising (Figure 3). A learning rate scheduler reducing the learning rate by a factor of 0.5 when validation loss plateaued for 5 consecutive epochs, and early stopping with a patience of 50 epochs.
All pre-training was conducted on the baskerville HPC cluster, on single a40 and a80 GPUs.
@@ 3.4 Fine-tuning
To fine-tune our pre-trained Perceiver-VAE, we freeze the encoder weights and train a lightweight classification head on top of the learned latent representations [6]. For a given input light curve, the frozen encoder produces a mean latent vector $\mu$, which is passed directly to the classification head. The optimisation objective is therefore simply the categorical cross-entropy loss:
$\mathcal{L}_{\text{fine-tune}}=\mathcal{L}_{\text{class}}=-\frac{1}{m}\sum_{i=1}^{m}\sum_{c=1}^{C}y_{ic}\log(\hat{y}_{ic}),$ (8)
across $m$ samples and $C$ classes, where $y_{ic}$ are the true labels and $\hat{y}_{ic}$ the predicted probabilities. This approach follows standard practice for fine-tuning foundation models, using the pre-trained encoder purely as a feature extractor.
The classification head is a two-layer MLP: a linear projection to a hidden dimension, followed by ReLU activation, dropout regularisation ($p=0.3$), and a final linear projection to the number of output classes.
We trained using the Adam optimiser with a learning rate of $1\times 10^{-3}$ for all fine-tuning conditions. Batch size and maximum epochs were set per task: anomaly detection used a batch size of 32 for up to 50 epochs, and motion prediction used a batch size of 64 for up to 150 epochs, reflecting the difference in dataset size between the two tasks. To address class imbalance in the motion prediction task (Table 6), we applied inverse-frequency class weighting to the cross-entropy loss. The best checkpoint was selected by validation ROC AUC.
To evaluate the contribution of pre-training, we additionally train classification heads on randomly initialised Perceiver encoders of equivalent scale, with weights drawn from the default PyTorch initialisation and held frozen throughout fine-tuning. These random encoder (r.e.) variants are otherwise identical to the pre-trained models, isolating the effect of pre-training on downstream performance.
3.4.1 Anomaly Detection Fine-tuning
For anomaly detection, the classification head produces a binary output distinguishing anomalous from normal light curves. We evaluated on a stratified held-out test set comprising 20% of the CASSANDRA dataset, reporting both accuracy and ROC AUC.
To evaluate label efficiency, we additionally conducted a $k$-shot study comparing the pre-trained Perceiver against the CNN baseline. For each value of $\in\{2,5,10,25\}, kk$ labelled examples per class were sampled from the training pool using independent random seeds across 20 independent runs. Each run trained the classification head (or CNN) for 50 epochs using the Adam optimiser with a learning rate of $1\times 10^{-3}$ and batch size of 32, and was evaluated on the same stratified held-out test set used in the main experiments.
3.4.2 Motion Prediction Fine-tuning
For motion prediction, the classification head produces a six-class output corresponding to the grouped motion classes described in Table 6. We similarly evaluated on a stratified 20% held-out test set, reporting per-class precision, recall, F1-score, and ROC AUC.
@@ 3.5 Synthetic Data Generation
Here we use the fine-tuned motion prediction classifier to generate a set of high confidence ’query’ curves. The algorithm first identifies some reference curves above a given confidence threshold (0.9) for a requested motion class e.g., tumbling. Then, it uses these to index the learned latent space of the model, and samples around that latent space by adding Gaussian noise to the highest activating neurons. For visual clarity in Figure 9, Gaussian smoothing ($\sigma=2.0$) is applied to the displayed synthetic curves to aid inspection of global morphological features. The GP validation (Section 2.7, Figure 10) is performed on unsmoothed outputs, confirming that quantitative synthetic curve quality is maintained without smoothing.
To quantitatively validate generated synthetic light curves, we fitted a Gaussian Process (GP) to the real light curve distribution for each motion class and evaluated whether synthetic curves fell within the learned distributional bounds.
For each class, high-confidence real curves (predicted class probability >>0.75 from the fine-tuned motion classifier) were selected, and a GP was fitted to these using an RBF kernel with additive white noise: $k(t,t^{\prime})=\text{RBF}(l{=}0.2)+\text{WhiteKernel}(\sigma^{2}{=}0.01)$, with kernel hyperparameters refined via three restarts of marginal likelihood optimisation and outputs normalised prior to fitting. This provides a non-parametric estimate of the mean and standard deviation of the real light curve distribution at each of the 128 resampled time steps.
Each synthetic curve was scored by computing the fraction of time steps for which the absolute deviation from the GP posterior mean fell within two standard deviations:
$s=\frac{1}{128}\sum_{i=1}^{128}\mathbf{1}\!\left[|\hat{y}_{i}-\mu_{i}|<2\sigma_{i}\right],$ (9)
where $\mu_{i}$ and $\sigma_{i}$ are the GP posterior mean and standard deviation at time step $i$, and $\hat{y}_{i}$ is the synthetic curve value. The class-level score is the mean of $s$ across all synthetic curves for that class. This treats the GP as a distributional reference for plausible real curves, such that a high score indicates synthetic curves are statistically consistent with real observations under the GP model.
3.6 Model Compression
To assess parameter efficiency of the large Perceiver-VAE, we applied global unstructured magnitude pruning [10] separately to the Perceiver layers and the VAE bottleneck (mean and log-variance projections), sweeping sparsity from 10% to 90%. Following pruning, batch normalisation statistics were recalibrated on 500 training samples prior to evaluation on the held-out test set. Post-training dynamic quantisation was applied to the full model, converting all linear layers to INT8 precision using PyTorch’s quantize_dynamic. No calibration data was required.
@@ 4 Outlook
Our work is the first to demonstrate that a VAE architecture can effectively learn meaningful representations from light curve data and be fine-tuned for multiple downstream tasks including detecting anomalous space object behaviour and motion prediction. We have also shown that our approach scales well to large datasets while maintaining computational efficiency, which is critical considering the rapidly growing amount of Space Object data. Here, we developed a two-stage approach, first through unsupervised pre-training, whereby the model learns rich features which aid reconstruction of real light curves. Then we performed supervised fine-tuning, which leverages those features for both anomaly classification, reaching 85% accuracy with a 0.92 ROC AUC score, and attitude mode motion classification, achieving 82.7% test accuracy, with 0.95 ROC AUC scores. Additionally, we demonstrate that the same architectural framework and approach can simultaneously achieve multiple space object behavioural analysis goals (i.e., anomaly detection, attitude mode classification, and synthetic data generation), provided that the representations learned during pre-training are diverse and rich, and that these are coupled with high-fidelity fine-tuning simulation data. Although in this study we identify several predicted anomalous light curve modes, some present clearly anomalous behaviour (e.g., rapid tumbling), while for others the underlying behaviour is less clear. A further direction is per-object anomaly characterisation: for well-represented objects in the pre-training data, fitting a Gaussian Process to the distribution of typical light curves for a specific object and comparing high reconstruction error curves against this baseline would allow more rigorous validation of whether these curves represent genuinely atypical behaviour for that object specifically. Systematic cataloguing and analysis of simulation conditions that lead to these various curve morphologies would allow us to verify whether particular erroneous space object dynamics produce these curves (e.g., in Fig 7C, F). A further valuable direction is a systematic pre-training scaling study, training models on increasingly large subsets of the MMT-9 dataset to characterise how pre-training data volume impacts performance. This is of particular relevance to the SDA context, where pre-training data may grow substantially as sensor networks expand, but detailed labels remain sparse.
@@
From this work, several key directions for further development emerge. Firstly, development of robust quantitative benchmarks for AI synthetic data. Whilst our approach encourages plausibility in the generated curves through reference-based sampling, systematic evaluation frameworks are needed to validate produced synthetic data beyond visual inspection/sense checks. Whilst we provide GP-based validation of synthetic light curve quality in this work, further quantitative evaluation metrics such as KL divergence, power spectral density comparisons, or Earth mover distance between real and synthetic distributions remain valuable directions for future work. For example, these frameworks could assess statistical distribution matching between real and AI generated datasets, and evaluate like-for-like comparisons in downstream tasks trained on real data, numerically simulated data, and/or data generated through our approach (i.e., generative AI data). Finally, physics-informed neural networks (PINNs) could be used to quantify physical plausibility. Briefly, these networks combine a data loss term with a physics loss term. For example, for light curves this might be a reflectivity model as outlined in [18]. At inference time, a well-trained PINN will give a good indicator as to whether a synthetically generated light curve is obeying the underlying physics due to the value of the physics loss term.
A key limitation of the current work is the simulation-to-real gap: both fine-tuning datasets (CASSANDRA for anomaly detection and GRIAL for motion prediction) are simulated, whilst pre-training data and any operational inference data are observations. Whilst the pre-trained model demonstrates some unsupervised anomaly discrimination on the CASSANDRA dataset (ROC AUC 0.71 without any labels), the domain gap between real and simulated curves limits threshold-based detection, motivating fine-tuning as a necessary bridging step. Closing this gap will require either real labelled anomaly data, which is currently scarce due to national security and commercial sensitivities, or domain adaptation techniques that better align simulated and real light curve distributions. Contrastive approaches using paired real and simulated observations of the same object could be a promising direction, as could physics-informed data augmentation to make simulated curves more representative of real observational conditions.
Finally, magnitude pruning experiments (Section 2.8) suggest the large model contains substantial redundancy in its Perceiver layers, motivating future work on more parameter-efficient pre-training strategies. Similarly, the negligible accuracy cost of INT8 quantisation suggests that deployment of large-scale space object characterisation models in resource-constrained operational environments is feasible.
@@
A substantial challenge direction for further development is multi-modal fusion - i.e., integrating the light curve analysis presented here with other observation types (e.g., hyperspectral data, orbits, radar etc.) Indeed, part of the decision to use a Perceiver based architecture was to readily enable extension to other data modalities, as it is well suited to this through its cross-attention and latent bottleneck design, which could enable multiple modalities to be reduced to a common set of embeddings. A natural extension of future work is to exploit the Perceiver’s ability to handle variable-length inputs directly, removing the uniform resampling constraint applied here and allowing the model to process light curves at their native sampling rates. However, there are substantial technical challenges in multimodal fusion, including developing an appropriate fusion strategy at different levels of the pre-training/fine-tuning pipeline. One such method to fuse information from other sensors with light curves could include implementation of contrastive learning using paired observations. For example, positive pairs (same space object at different dates) and negative pairs (different objects on the same date) would help develop more discriminative features for space object characterisation. Additionally incorporating metadata prediction tasks as part of the self-supervised framework could help capture characteristic signatures of specific satellites, or debris. For example, a fine-tuned metadata prediction FM could generate a likely light curve given an orbital eccentricity, which could help unpick the relationships between satellite structure/behaviour and orbital path.
Our initial framework provides promising results, and as the amount of data available increases, a basis for further analysis. We suggest Foundation Models, in particular those that exploit multimodal datasets, could become powerful tools for automated space object monitoring and space situational awareness. As orbital populations continue to grow, automated approaches such as ours will become increasingly crucial for maintaining space safety & sustainability.
@@
5 Data Availability Statement
Research data are not shared.
Supplementary: Baseline Model Architectures
All baseline models take a light curve of length 128 as input and are trained from scratch using the Adam optimiser with a learning rate of $1\times 10^{-3}$, batch size of 32, and dropout of $p=0.3$ throughout.
RNN: A two-layer recurrent neural network with tanh activation and hidden size 64, following a standard stacked configuration to capture both short and long-range sequential dependencies. The final hidden state is passed to a linear output projection.
LSTM: A two-layer Long Short-Term Memory network with hidden size 64, again using a stacked configuration for multi-scale temporal modelling. The final hidden state is passed to a linear output projection.
CNN: Three successive 1D convolutional blocks with progressively increasing channel depths (32, 64, 128) and kernel sizes of 7, 5, and 3 respectively, to capture hierarchical temporal features at decreasing scales. Each block is followed by ReLU activation and max pooling. Global average pooling precedes a linear output projection.
Transformer: A two-layer Transformer encoder with model dimension $d_{\text{model}}$=64, 4 attention heads, and feedforward dimension 256, following the standard 1:4 model-to-feedforward scaling convention [27]. Input timesteps are projected to $d_{\text{model}}$ before encoding, and global average pooling over the sequence precedes a linear output projection.